OpenStudy (clara1223):

Determine the values of the constants B and C so that the function given below is differentiable. f(x)={8x^3 when x≤1; (Bx^2)+Cx when x>1} a) {B=−24,C=24} b) {B=24,C=32} c) {B=16,C=−8} d) {B=24,C=40} e) {B=−48,C=−8}

OpenStudy (anonymous):

plug in \(x=1\) set them equal take the derivative, plug in \(x=1\) and set those equal solve for B, C

OpenStudy (anonymous):

for example the first equation you get is \[8=B+C\]

OpenStudy (clara1223):

once i have 8=B+C and 24=2B+C how do I solve for B and C?

OpenStudy (clara1223):

@satellite73

OpenStudy (anonymous):

same way you solve any system of equations \[B+C=8\\ 2B+C=24\] elimination or substitution

OpenStudy (anonymous):

or just subtract equation one from equation two and get \[B=16\] right away

OpenStudy (clara1223):

figured it out! thanks!